# Can we determine the exact size of the nucleon?: A comprehensive study of different radii

### v1

@@ I Introduction
Considering the wide range of values obtained for these radii from various experimental, theoretical, and phenomenological studies, the question arises: Can we interpret these radii as the size of the nucleon? It appears that interpreting nucleon radii as the “size” of the nucleon is a nuanced and context-dependent topic. While these radii provide measures of spatial distributions associated with different properties of the nucleon-such as charge, magnetism, weak interactions, mechanical stress, and mass distribution-they do not all correspond to a single, unified concept of size. To be more precise, each radius offers a measure of the nucleon’s spatial extent in the context of a specific physical property, but there is no single, universally agreed-upon definition of the nucleon’s size. For example, the charge radius is often interpreted as the size of the nucleon in terms of its electric charge distribution Pohl et al. (2010); Karr et al. (2020), but this interpretation is open to criticism given the multifaceted nature of nucleon structure and the differences between various radii Kaiser and Weise (2024); Petrov (2024). Thus, while each radius provides valuable insights into the nucleon’s internal structure, they collectively highlight the complexity of defining a single “size” for the nucleon.
@@ II The charge and magnetic radii
where denotes the proton and neutron, respectively, and is the magnetic moment of the nucleon. While electron scattering provides precise measurements of the proton’s charge and magnetic radii, its application to the neutron is more challenging due to the neutron’s lack of a net electric charge. For the neutron, the measurements of the double polarization observables in quasi-elastic electron scattering from polarized deuterium or 3He targets is often used to determine the neutron electric FF, , though this method introduces additional theoretical uncertainties related to nuclear corrections and was not able to access at a sufficiently low range Atac et al. (2021a).
@@
Besides the direct experimental extractions of the nucleon charge and magnetic radii, there have been various analyses of the elastic scattering data Ridwan and Mart (2023); Lorenz et al. (2012); Lorenz and Meißner (2014); Kraus et al. (2014); Higinbotham et al. (2018); Zahra et al. (2022); Pacetti and Tomasi-Gustafsson (2021); Cui et al. (2021a); Sick and Trautmann (2017); Horbatsch and Hessels (2016); Lin et al. (2021a); Horbatsch (2020); Hagelstein and Pascalutsa (2019); Borisyuk and Kobushkin (2020); Borah et al. (2020); Filin et al. (2020, 2021); Barcus et al. (2020); Vaziri and Reza Shojaei (2023); Griffioen et al. (2016); Alarcón et al. (2019); Hill and Paz (2010); Zhou et al. (2019); Cui et al. (2021b); Sick (2018); Higinbotham et al. (2016); Belushkin et al. (2007); Borisyuk (2010); Lorenz et al. (2015); Graczyk and Juszczak (2014); Alarcón et al. (2020); Lee et al. (2015); Arrington and Sick (2015); Horbatsch et al. (2017); Lin et al. (2021b, 2022); Goharipour et al. (2024); Atac et al. (2021b); Gramolin and Russell (2022); Hayward and Griffioen (2020); Paz (2021); Atoui et al. (2024); Albloushi et al. (2024) aimed at resolving the discrepancies observed between different experimental techniques. These studies employ diverse methodologies, including conformal mapping, dispersive approaches, Bayesian frameworks, and statistical sampling techniques, to extract the nucleon radii, particularly the proton charge radius, from experimental data Gao and Vanderhaeghen (2022). Most analyses rely on high-precision data from the Mainz A1 experiment Bernauer et al. (2010, 2014) and, more recently, the PRad experiment Xiong et al. (2019). These datasets provide a robust foundation for extracting the proton charge radius with reduced uncertainties. Many analyses, such as those by Lorenz et al. Lorenz et al. (2012); Lorenz and Meißner (2014); Lorenz et al. (2015), Griffioen et al. Griffioen et al. (2016), Alarcón et al. Alarcón et al. (2019, 2020), and Lin et al. Lin et al. (2021b, 2022), yield values consistent with the smaller muonic hydrogen results ( fm). These studies often incorporate advanced theoretical constraints, such as dispersive improvements, chiral perturbation theory (PT), and Two-Photon-Exchange (TPE) corrections, to ensure precision. On the other hand, many analyses, such as those by Hill and Paz Hill and Paz (2010), Graczyk and Juszczak Graczyk and Juszczak (2014), Zhou et al. Zhou et al. (2019), Cui et al. Cui et al. (2021b, a), and Paz Paz (2021), employ model-independent techniques (e.g., conformal mapping, Bayesian methods, or the Schlessinger Point Method) to minimize biases from assumed functional forms. These analyses also often lead to values consistent with the smaller muonic hydrogen results. However, some studies, such as Lee et al. Lee et al. (2015), Sick Sick (2018), Arrington and Sick Arrington and Sick (2015), and Gramolin and Russell Gramolin and Russell (2022), report larger values, closer to older electron scattering results.
@@
The differences between these analyses often stem from variations in methodology, data selection, and treatment of systematic uncertainties. For example, some studies, like Griffioen et al. Griffioen et al. (2016), inflate systematic uncertainties to account for normalization issues, while others, like Alarcón et al. Alarcón et al. (2019, 2020), incorporate theoretical uncertainties from spectral functions or PT. As another example, Sick and Trautmann Sick and Trautmann (2017) argued that neglecting higher moments in some analyses leads to smaller radius values, and Horbatsch et al. Horbatsch et al. (2017) addressed this by fixing higher moments using PT. It is also worth noting in this context the following analyses. Horbatsch Horbatsch (2020) analyzed the PRad data by taking the logarithm to yield a -dependent radius function. Atac et al. Atac et al. (2021b) extracted both the proton and neutron charge radii from a global analysis of the proton and neutron elastic form factors by carrying out a flavor decomposition of these FFs under charge symmetry in the light-cone frame formulation. Gramolin and Russell Gramolin and Russell (2022) proposed a novel method for extracting , which relates the radius of the proton to its transverse charge density, the two-dimensional Fourier transform of the Dirac form factor. Filin et al. Filin et al. (2020, 2021) extracted the neutron charge radius through a comprehensive analysis of the deuteron charge and quadrupole form factor data in chiral effective field theory. Atoui et al. Atoui et al. (2024) determined through a global analysis of proton electric form factor data based on an integral method. The MMGPDs Collaboration Goharipour et al. (2024) determined the charge and magnetic radii of both the proton and neutron through a simulation analysis of a wide range of elastic scattering data, considering the relation between the Pauli and Dirac FFs (and hence the Sachs FFs) and the generalized parton distributions (GPDs) at zero skewness. Finally, note that the extracted radius values can be sensitive to the range of the data included, as mentioned before. For example, Ridwan and Mart Ridwan and Mart (2023), Higinbotham et al. Higinbotham et al. (2016), Zhou et al. Zhou et al. (2019), and Pacetti and Tomasi-Gustafsson Pacetti and Tomasi-Gustafsson (2021) highlight the impact of data selection on the final results. The values extracted by Zhou et al. Zhou et al. (2019) for the proton charge radius range from 0.83 to 0.88 fm, while the mixed result is fm. As another example, Hill and Paz Hill and Paz (2010) indicated that adding the neutron and data in the analysis leads to a decrease in the uncertainty of the proton charge radius. In addition, Barcus et al. Barcus et al. (2020) indicated how analytic choices, such as the function that is being used or even the binning of the data, can affect the extraction of electromagnetic FFs from elastic electron scattering cross Section data and hence the extracted proton charge radius.
@@
In addition to experimental measurements and phenomenological analyses, a variety of theoretical approaches have been developed to calculate the electromagnetic FFs or the charge and magnetic radii of the nucleon Ramalho and Tsushima (2011); Trinhammer (2021); Shanahan et al. (2014a, b); Chambers et al. (2017); Alexandrou et al. (2017a, 2020, 2019); Hasan et al. (2018); Park et al. (2022); Ishikawa et al. (2021); Shintani et al. (2019); Tsuji et al. (2024); Jang et al. (2020a); Djukanovic et al. (2021, 2024a, 2024b); Syritsyn et al. (2023); Abidin and Carlson (2009); Brodsky et al. (2015); Sufian et al. (2017); Gutsche et al. (2020); Contreras et al. (2021); Mamo and Zahed (2023); Ahmady et al. (2021); Xu et al. (2021); Mondal et al. (2020); Maris and Roberts (2003); Eichmann et al. (2010, 2016); Burkert and Roberts (2019); Ding et al. (2023); Ferreira and Papavassiliou (2023); Barabanov et al. (2021); Yao et al. (2024a); Gasser et al. (1988); Kubis and Meissner (2001); Alarcon et al. (2014); Bauer et al. (2012); Liu et al. (2014); Flores-Mendieta and Rivera-Ruiz (2015); He and Wang (2018); Hiller Blin (2017); Yang and Wang (2020); Bar and Colic (2021); Alvarado et al. (2023); Aung et al. (2024); Berger et al. (2004); Julia-Diaz et al. (2004); Sharma and Dahiya (2013); Plessas (2015); Balitsky et al. (1985); Lenz et al. (2004); Castillo et al. (2005); Braun et al. (2006); Aliev et al. (2008); Anikin et al. (2013); Brodsky et al. (2022); Gao et al. (2023); Xing et al. (2023); Ramalho and Peña (2024); Faessler et al. (2010); Kuzmin et al. (2025). These approaches are rooted in QCD, but they employ different strategies to address the non-perturbative nature of QCD at low energies. One of the most prominent methods is lattice QCD, which discretizes spacetime and numerically solves QCD on a finite grid. Lattice QCD has made significant progress in recent years, providing first-principles calculations of nucleon FFs and radii with increasingly precise results Shanahan et al. (2014a, b); Chambers et al. (2017); Hasan et al. (2018); Park et al. (2022); Alexandrou et al. (2017a, 2019, 2020); Shintani et al. (2019); Ishikawa et al. (2021); Tsuji et al. (2024); Jang et al. (2020a); Djukanovic et al. (2021, 2024a, 2024b); Syritsyn et al. (2023). Another important framework is holographic QCD which is a theoretical framework that uses the anti de Sitter space/conformal field theory (AdS/CFT) correspondence to study QCD in the non-perturbative regime Abidin and Carlson (2009); Brodsky et al. (2015); Sufian et al. (2017); Gutsche et al. (2020); Contreras et al. (2021); Mamo and Zahed (2023); Ahmady et al. (2021); Mondal et al. (2020); Xu et al. (2021). Additionally, Dyson-Schwinger equations (DSEs) and Bethe-Salpeter approaches offer a continuum formulation of QCD, enabling the calculation of form factors by solving integral equations for quark and gluon propagators Maris and Roberts (2003); Eichmann et al. (2010, 2016); Burkert and Roberts (2019); Ding et al. (2023); Ferreira and Papavassiliou (2023); Barabanov et al. (2021); Yao et al. (2024a). It is worth pointing in this context to various chiral theories and models Gasser et al. (1988); Kubis and Meissner (2001); Alarcon et al. (2014); Bauer et al. (2012); Liu et al. (2014); Flores-Mendieta and Rivera-Ruiz (2015); He and Wang (2018); Hiller Blin (2017); Yang and Wang (2020); Bar and Colic (2021); Alvarado et al. (2023); Aung et al. (2024). For example, PT exploits the symmetries of QCD in the low-energy regime and provides systematic expansions in terms of the pion mass and momentum. PT has been particularly useful for understanding the pion cloud contributions to nucleon FFs and radii. There are also other approaches such as the constituent quark models, which describe nucleons as bound states of three quarks interacting via phenomenological potentials Berger et al. (2004); Julia-Diaz et al. (2004); Sharma and Dahiya (2013); Plessas (2015) and LCSR, which is a theoretical framework used to study the properties of hadrons by combining the principles of QCD sum rules with the light-cone distribution amplitudes of hadrons Balitsky et al. (1985); Lenz et al. (2004); Castillo et al. (2005); Braun et al. (2006); Aliev et al. (2008); Anikin et al. (2013). These theoretical frameworks, while differing in their assumptions and methodologies, collectively provide valuable insights into the electromagnetic structure of the nucleon and help bridge the gap between experimental data and fundamental QCD principles.
@@
Figure 1 (left panel) shows a comparison between the results obtained for the proton charge radius from different experimental techniques. In this figure (as well as the next figure), the right band and the left wider band correspond to the CODATA 2014 Mohr et al. (2016) and CODATA 2018 Tiesinga et al. (2021) values, respectively, while the left narrow band represents the PDG 2024 Navas et al. (2024) value. For studies where multiple sources of uncertainties were provided, we have added them in quadrature to obtain a single combined error. As can be seen, the results clearly illustrate the proton charge radius puzzle. While muonic hydrogen spectroscopy (diamonds) consistently yields smaller values of Pohl et al. (2010); Antognini et al. (2013), the results from ordinary hydrogen spectroscopy (circles) span a wide range. Notably, two recent measurements from hydrogen spectroscopy Grinin et al. (2020); Brandt et al. (2022) have reported moderate values for . Among the scattering experiments Xiong et al. (2019); Bernauer et al. (2010); Zhan et al. (2011); Mihovilovič et al. (2021) (squares), only the PRad experiment Xiong et al. (2019) is in agreement with the muonic hydrogen spectroscopy results. It is also worth noting that muonic hydrogen spectroscopy achieves significantly smaller uncertainties compared to both scattering experiments and ordinary hydrogen spectroscopy. Overall, these experimental results cover the interval fm. A simple combination of the scattering experiments yields fm, while the corresponding value from hydrogen spectroscopy is fm, which is approximately 1.85% smaller than the average of the scattering experiment results. The simple combination of all experimental results leads to fm.
@@
In the right panel of Fig. 1, we compare the results obtained for from various theoretical approaches. The primary results come from lattice QCD Park et al. (2022); Alexandrou et al. (2019, 2020); Shintani et al. (2019); Ishikawa et al. (2021); Tsuji et al. (2024); Jang et al. (2020a); Djukanovic et al. (2021, 2024a) (circles). Additionally, we include results from holographic QCD Mamo and Zahed (2023); Ahmady et al. (2021); Xu et al. (2021) (stars) and chiral calculations Liu et al. (2014); Yang and Wang (2020) (triangles). As can be seen, most theoretical results exhibit significant uncertainties. Notably, lattice QCD predictions tend to yield smaller values for compared to other theoretical approaches and experimental results. Although the NME Collaboration Park et al. (2022) obtained a relatively large value, its uncertainty is also substantial, with the lower band even overlapping the lower range of all experimental results. Overall, the theoretical results span the intervals fm (including lattice results) and fm (excluding lattice results). The corresponding simple combinations are fm and fm, respectively. The simple combination of the lattice results alone yields fm, which is remarkably small. Based on these results, it is evident that theoretical approaches generally predict smaller values for compared to experimental measurements.
@@
Figure 2 shows a comparison between the results obtained for the proton charge radius from different analyses and phenomenological studies Hill and Paz (2010); Lorenz et al. (2012, 2015); Griffioen et al. (2016); Alarcón et al. (2019, 2020); Cui et al. (2021b); Lee et al. (2015); Sick (2018); Arrington and Sick (2015); Horbatsch et al. (2017); Higinbotham et al. (2016); Lin et al. (2021b, 2022); Zhou et al. (2019); Atac et al. (2021b); Gramolin and Russell (2022); Hayward and Griffioen (2020); Paz (2021); Borisyuk (2010); Graczyk and Juszczak (2014); Atoui et al. (2024); Albloushi et al. (2024); Belushkin et al. (2007); Goharipour et al. (2024). Note that for Ref. Lee et al. (2015), we have taken the value obtained by a simple combination of the Mainz and world values. Overall, the results are predominantly located in the region corresponding to smaller values of . The interval covered by these studies is fm. A simple combination of all results leads to fm. Interestingly, this value is in excellent agreement with the simple combination of all experimental results, which is fm. This is also consistent with the recent findings of the MMGPDs Collaboration Goharipour et al. (2024), which demonstrated that analyzing all available electron-nucleon scattering data simultaneously within the framework of GPDs yields an average value of fm, albeit with a relatively large uncertainty. It is also insightful to combine all results obtained from experimental, theoretical, and phenomenological efforts. Doing so, we obtain fm.
@@
In the right panel, we compare the results obtained from different analyses and phenomenological studies Lorenz et al. (2012); Lorenz and Meißner (2014); Lorenz et al. (2015); Alarcón et al. (2020); Lee et al. (2015); Arrington and Sick (2015); Horbatsch et al. (2017); Lin et al. (2021b, 2022); Cui et al. (2021a); Borah et al. (2020); Borisyuk (2010); Graczyk and Juszczak (2014); Belushkin et al. (2007); Goharipour et al. (2024). These results cover the interval fm, with a simple combination of fm. This value is smaller than the combined experimental results but larger than the combined theoretical results. It is also worth noting that the difference between the combined experimental and phenomenological results (0.016 fm) is more significant compared to the case of . Finally, combining all experimental, theoretical, and phenomenological results, we obtain fm.

### v3

@@ I Introduction
Considering the wide range of values obtained for these radii from various experimental, theoretical, and phenomenological studies, the question arises: Can we interpret these radii as the size of the nucleon? It appears that interpreting nucleon radii as the “size” of the nucleon is a nuanced and context-dependent topic. While these radii provide measures of spatial distributions associated with different properties of the nucleon-such as charge, magnetism, weak interactions, mechanical stress, and mass distribution-they do not all correspond to a single, unified concept of size. To be more precise, each radius offers a measure of the nucleon’s spatial extent in the context of a specific physical property, but there is no single, universally agreed-upon definition of the nucleon’s size. For example, the charge radius is often interpreted as the size of the nucleon in terms of its electric charge distribution Pohl et al. (2010); Karr et al. (2020), but this interpretation is open to criticism given the multifaceted nature of nucleon structure and the differences between various radii Liu et al. (1999); Kaiser and Weise (2024); Petrov (2024). Thus, while each radius provides valuable insights into the nucleon’s internal structure, they collectively highlight the complexity of defining a single “size” for the nucleon.
@@ II The charge and magnetic radii
where denotes the proton and neutron, respectively, and is the magnetic moment of the nucleon. While electron scattering provides precise measurements of the proton’s charge and magnetic radii, its application to the neutron is more challenging due to the neutron’s lack of a net electric charge. For the neutron, the measurements of the double polarization observables in quasi-elastic electron scattering from polarized deuterium or 3He targets is often used to determine the neutron electric FF, , though this method introduces additional theoretical uncertainties related to nuclear corrections and was not able to access at a sufficiently low range Atac et al. (2021a). It should be also noted that radii are not fundamental quantities in quantum field theory (QFT). The conventional definition of in Eq. (1), rooted in non-relativistic form factor expansions at low , is inadequate for relativistic systems Miller (2019, 2007); Chen and Lorcé (2022, 2023). A more rigorous framework is provided by impact-parameter-dependent GPDs Ji (1998); Burkardt (2000, 2003); Singireddy et al. (2025), which avoid the ambiguities of a universal radius definition. These QFT-based approaches highlight the limitations of interpreting as a strict measure of spatial extent.
@@
Besides the direct experimental extractions of the nucleon charge and magnetic radii, there have been various analyses of the elastic scattering data Ridwan and Mart (2023); Lorenz et al. (2012); Kelly (2002); Lorenz and Meißner (2014); Kraus et al. (2014); Higinbotham et al. (2018); Zahra et al. (2022); Pacetti and Tomasi-Gustafsson (2021); Cui et al. (2021a); Sick and Trautmann (2017); Horbatsch and Hessels (2016); Lin et al. (2021a); Horbatsch (2020); Hagelstein and Pascalutsa (2019); Borisyuk and Kobushkin (2020); Borah et al. (2020); Filin et al. (2020, 2021); Barcus et al. (2020); Vaziri and Reza Shojaei (2023); Griffioen et al. (2016); Alarcón et al. (2019); Hill and Paz (2010); Zhou et al. (2019); Cui et al. (2021b); Sick (2018); Higinbotham et al. (2016); Belushkin et al. (2007); Borisyuk (2010); Lorenz et al. (2015); Graczyk and Juszczak (2014); Alarcón et al. (2020); Lee et al. (2015); Arrington and Sick (2015); Horbatsch et al. (2017); Lin et al. (2021b, 2022); Goharipour et al. (2025a); Atac et al. (2021b); Gramolin and Russell (2022); Hayward and Griffioen (2020); Paz (2021); Atoui et al. (2024); Albloushi et al. (2024) aimed at resolving the discrepancies observed between different experimental techniques. These studies employ diverse methodologies, including conformal mapping, dispersive approaches, Bayesian frameworks, and statistical sampling techniques, to extract the nucleon radii, particularly the proton charge radius, from experimental data Gao and Vanderhaeghen (2022). Most analyses rely on high-precision data from the Mainz A1 experiment Bernauer et al. (2010, 2014) and, more recently, the PRad experiment Xiong et al. (2019). These datasets provide a robust foundation for extracting the proton charge radius with reduced uncertainties. Many analyses, such as those by Lorenz et al. Lorenz et al. (2012); Lorenz and Meißner (2014); Lorenz et al. (2015), Griffioen et al. Griffioen et al. (2016), Alarcón et al. Alarcón et al. (2019, 2020), and Lin et al. Lin et al. (2021b, 2022), yield values consistent with the smaller muonic hydrogen results ( fm). These studies often incorporate advanced theoretical constraints, such as dispersive improvements, chiral perturbation theory (PT), and Two-Photon-Exchange (TPE) corrections, to ensure precision. On the other hand, many analyses, such as those by Hill and Paz Hill and Paz (2010), Graczyk and Juszczak Graczyk and Juszczak (2014), Zhou et al. Zhou et al. (2019), Cui et al. Cui et al. (2021b, a), and Paz Paz (2021), employ model-independent techniques (e.g., conformal mapping, Bayesian methods, or the Schlessinger Point Method) to minimize biases from assumed functional forms. These analyses also often lead to values consistent with the smaller muonic hydrogen results. However, some studies, such as Lee et al. Lee et al. (2015), Sick Sick (2018), Arrington and Sick Arrington and Sick (2015), and Gramolin and Russell Gramolin and Russell (2022), report larger values, closer to older electron scattering results.
@@
The differences between these analyses often stem from variations in methodology, data selection, and treatment of systematic uncertainties. For example, some studies, like Griffioen et al. Griffioen et al. (2016), inflate systematic uncertainties to account for normalization issues, while others, like Alarcón et al. Alarcón et al. (2019, 2020), incorporate theoretical uncertainties from spectral functions or PT. As another example, Sick and Trautmann Sick and Trautmann (2017) argued that neglecting higher moments of the charge density distribution in some analyses leads to smaller radius values, and Horbatsch et al. Horbatsch et al. (2017) addressed this by fixing higher moments using PT. It is also worth noting in this context the following analyses. Horbatsch Horbatsch (2020) analyzed the PRad data by taking the logarithm to yield a -dependent radius function. Atac et al. Atac et al. (2021b) extracted both the proton and neutron charge radii from a global analysis of the proton and neutron elastic form factors by carrying out a flavor decomposition of these FFs under charge symmetry in the light-cone frame formulation. Gramolin and Russell Gramolin and Russell (2022) proposed a novel method for extracting , which relates the radius of the proton to its transverse charge density, the two-dimensional Fourier transform of the Dirac form factor. Filin et al. Filin et al. (2020, 2021) extracted the neutron charge radius through a comprehensive analysis of the deuteron charge and quadrupole form factor data in chiral effective field theory. Atoui et al. Atoui et al. (2024) determined through a global analysis of proton electric form factor data based on an integral method.
The MMGPDs Collaboration Goharipour et al. (2025a) determined the charge and magnetic radii of both the proton and neutron through a simulation analysis of a wide range of elastic scattering data, considering the relation between the Pauli and Dirac FFs (and hence the Sachs FFs) and the generalized parton distributions (GPDs) at zero skewness. While GPDs offer a powerful framework for investigating nucleon structure through their unified description of spatial and momentum-space distributions, their application to radius extraction may need further consideration. Actually, the integration over the momentum fraction may introduce additional model dependence compared to direct form factor analyses, potentially affecting the precision of radius determinations. Nevertheless, note that the extracted GPDs leads to consistent results for other observables and quantities such as the electromagnetic, axial, gravitational, and transition FFs Hashamipour et al. (2023); Irani et al. (2023); Goharipour et al. (2024). Finally, note that the extracted radius values can be sensitive to the range of the data included, as mentioned before. For example, Ridwan and Mart Ridwan and Mart (2023), Higinbotham et al. Higinbotham et al. (2016), Zhou et al. Zhou et al. (2019), and Pacetti and Tomasi-Gustafsson Pacetti and Tomasi-Gustafsson (2021) highlight the impact of data selection on the final results. The values extracted by Zhou et al. Zhou et al. (2019) for the proton charge radius range from 0.83 to 0.88 fm, while the mixed result is fm. As another example, Hill and Paz Hill and Paz (2010) indicated that adding the neutron and data in the analysis leads to a decrease in the uncertainty of the proton charge radius. In addition, Barcus et al. Barcus et al. (2020) indicated how analytic choices, such as the function that is being used or even the binning of the data, can affect the extraction of electromagnetic FFs from elastic electron scattering cross section data and hence the extracted proton charge radius.
@@
In addition to experimental measurements and phenomenological analyses, a variety of theoretical approaches have been developed to calculate the electromagnetic FFs or the charge and magnetic radii of the nucleon Yennie et al. (1957); Ramalho and Tsushima (2011); Trinhammer (2021); Shanahan et al. (2014a, b); Chambers et al. (2017); Alexandrou et al. (2017a, 2020, 2019); Hasan et al. (2018); Park et al. (2022); Ishikawa et al. (2021); Shintani et al. (2019); Tsuji et al. (2024); Jang et al. (2020a); Djukanovic et al. (2021, 2024a, 2024b); Syritsyn et al. (2023); Abidin and Carlson (2009); Brodsky et al. (2015); Sufian et al. (2017); Gutsche et al. (2020); Contreras et al. (2021); Mamo and Zahed (2023); Ahmady et al. (2021); Xu et al. (2021); Mondal et al. (2020); Maris and Roberts (2003); Eichmann et al. (2010, 2016); Burkert and Roberts (2019); Ding et al. (2023); Ferreira and Papavassiliou (2023); Barabanov et al. (2021); Yao et al. (2024a); Gasser et al. (1988); Kubis and Meissner (2001); Alarcon et al. (2014); Bauer et al. (2012); Liu et al. (2014); Flores-Mendieta and Rivera-Ruiz (2015); He and Wang (2018); Hiller Blin (2017); Yang and Wang (2020); Bar and Colic (2021); Alvarado et al. (2023); Aung et al. (2024); Berger et al. (2004); Julia-Diaz et al. (2004); Sharma and Dahiya (2013); Plessas (2015); Balitsky et al. (1985); Lenz et al. (2004); Castillo et al. (2005); Braun et al. (2006); Aliev et al. (2008); Anikin et al. (2013); Brodsky et al. (2022); Gao et al. (2023); Xing et al. (2023); Ramalho and Peña (2024); Faessler et al. (2010); Kuzmin et al. (2025). These approaches are rooted in QCD, but they employ different strategies to address the non-perturbative nature of QCD at low energies. One of the most prominent methods is lattice QCD, which discretizes spacetime and numerically solves QCD on a finite grid. Lattice QCD has made significant progress in recent years, providing first-principles calculations of nucleon FFs and radii with increasingly precise results Shanahan et al. (2014a, b); Chambers et al. (2017); Hasan et al. (2018); Park et al. (2022); Alexandrou et al. (2017a, 2019, 2020); Shintani et al. (2019); Ishikawa et al. (2021); Tsuji et al. (2024); Jang et al. (2020a); Djukanovic et al. (2021, 2024a, 2024b); Syritsyn et al. (2023). Another important framework is holographic QCD which is a theoretical framework that uses the anti de Sitter space/conformal field theory (AdS/CFT) correspondence to study QCD in the non-perturbative regime Abidin and Carlson (2009); Brodsky et al. (2015); Sufian et al. (2017); Gutsche et al. (2020); Contreras et al. (2021); Mamo and Zahed (2023); Ahmady et al. (2021); Mondal et al. (2020); Xu et al. (2021). Additionally, Dyson-Schwinger equations (DSEs) and Bethe-Salpeter approaches offer a continuum formulation of QCD, enabling the calculation of form factors by solving integral equations for quark and gluon propagators Maris and Roberts (2003); Eichmann et al. (2010, 2016); Burkert and Roberts (2019); Ding et al. (2023); Ferreira and Papavassiliou (2023); Barabanov et al. (2021); Yao et al. (2024a). It is worth pointing in this context to various chiral theories and models Gasser et al. (1988); Kubis and Meissner (2001); Alarcon et al. (2014); Bauer et al. (2012); Liu et al. (2014); Flores-Mendieta and Rivera-Ruiz (2015); He and Wang (2018); Hiller Blin (2017); Yang and Wang (2020); Bar and Colic (2021); Alvarado et al. (2023); Aung et al. (2024). For example, PT exploits the symmetries of QCD in the low-energy regime and provides systematic expansions in terms of the pion mass and momentum. PT has been particularly useful for understanding the pion cloud contributions to nucleon FFs and radii. There are also other approaches such as the constituent quark models, which describe nucleons as bound states of three quarks interacting via phenomenological potentials Berger et al. (2004); Julia-Diaz et al. (2004); Sharma and Dahiya (2013); Plessas (2015) and LCSR, which is a theoretical framework used to study the properties of hadrons by combining the principles of QCD sum rules with the light-cone distribution amplitudes of hadrons Balitsky et al. (1985); Lenz et al. (2004); Castillo et al. (2005); Braun et al. (2006); Aliev et al. (2008); Anikin et al. (2013). These theoretical frameworks, while differing in their assumptions and methodologies, collectively provide valuable insights into the electromagnetic structure of the nucleon and help bridge the gap between experimental data and fundamental QCD principles.
@@
Figure 1 (left panel) shows a comparison between the results obtained for the proton charge radius from different experimental techniques. In this figure (as well as the next figure), the right band and the left wider band correspond to the CODATA 2014 Mohr et al. (2016) and CODATA 2018 Tiesinga et al. (2021) values, respectively, while the left narrow band represents the PDG 2024 Navas et al. (2024) value. For studies where multiple sources of uncertainties were provided, we have added them in quadrature to obtain a single combined error. As can be seen, the results clearly illustrate the proton charge radius puzzle. While muonic hydrogen spectroscopy (diamonds) consistently yields smaller values of Pohl et al. (2010); Antognini et al. (2013), the results from ordinary hydrogen spectroscopy (circles) span a wide range. Notably, two recent measurements from hydrogen spectroscopy Grinin et al. (2020); Brandt et al. (2022) have reported moderate values for . Among the scattering experiments Xiong et al. (2019); Bernauer et al. (2010); Zhan et al. (2011); Mihovilovič et al. (2021) (squares), only the PRad experiment Xiong et al. (2019) is in agreement with the muonic hydrogen spectroscopy results. It is also worth noting that muonic hydrogen spectroscopy achieves significantly smaller uncertainties compared to both scattering experiments and ordinary hydrogen spectroscopy. Overall, these experimental results cover the interval fm. A simple combination of the scattering experiments yields fm, while the corresponding value from hydrogen spectroscopy is fm, which is approximately 1.85% smaller than the average of the scattering experiment results. The simple combination of all experimental results leads to fm. We note that these combined values represent unweighted averages of independent results, with uncertainties propagated through standard error analysis assuming no correlations. This approach is adopted throughout this article to calculate the combined values.
@@
In the right panel of Fig. 1, we compare the results obtained for from various theoretical approaches. The primary results come from lattice QCD Park et al. (2022); Alexandrou et al. (2019, 2020); Shintani et al. (2019); Ishikawa et al. (2021); Tsuji et al. (2024); Jang et al. (2020a); Djukanovic et al. (2021, 2024a) (circles). Additionally, we include results from holographic QCD Mamo and Zahed (2023); Ahmady et al. (2021); Xu et al. (2021) (stars) and chiral calculations Liu et al. (2014); Yang and Wang (2020) (triangles). As can be seen, most theoretical results exhibit significant uncertainties. The results clearly show that lattice QCD predictions have not yet converged. This variability complicates direct comparison with experimental determinations and highlights the need for continued improvements in lattice methodologies. For instance, although the NME Collaboration Park et al. (2022) obtained a relatively large value, its uncertainty is also substantial, with the lower band even overlapping the lower range of all experimental results. Overall, the theoretical results span the intervals fm (including lattice results) and fm (excluding lattice results). The corresponding simple combinations are fm and fm, respectively. The simple combination of the lattice results alone yields fm, indicating that lattice QCD predictions tend to yield smaller values for compared to other theoretical approaches and experimental results. Moreover, it is evident that theoretical approaches generally predict smaller values for compared to experimental measurements.
@@
Figure 2 shows a comparison between the results obtained for the proton charge radius from different analyses and phenomenological studies Hill and Paz (2010); Lorenz et al. (2012, 2015); Griffioen et al. (2016); Alarcón et al. (2019, 2020); Cui et al. (2021b); Lee et al. (2015); Sick (2018); Arrington and Sick (2015); Horbatsch et al. (2017); Higinbotham et al. (2016); Lin et al. (2021b, 2022); Zhou et al. (2019); Atac et al. (2021b); Gramolin and Russell (2022); Hayward and Griffioen (2020); Paz (2021); Borisyuk (2010); Graczyk and Juszczak (2014); Atoui et al. (2024); Albloushi et al. (2024); Belushkin et al. (2007); Goharipour et al. (2025a). Note that for Ref. Lee et al. (2015), we have taken the value obtained by a simple combination of the Mainz and world values. Overall, the results are predominantly located in the region corresponding to smaller values of . The interval covered by these studies is fm. A simple combination of all results leads to fm. Interestingly, this value is in excellent agreement with the simple combination of all experimental results, which is fm. This is also consistent with the recent findings of the MMGPDs Collaboration Goharipour et al. (2025a), which demonstrated that analyzing all available electron-nucleon scattering data simultaneously within the framework of GPDs yields an average value of fm, albeit with a relatively large uncertainty. It is also insightful to combine all results obtained from experimental, theoretical, and phenomenological efforts. Doing so, we obtain fm.
@@
In the right panel, we compare the results obtained from different analyses and phenomenological studies Lorenz et al. (2012); Lorenz and Meißner (2014); Lorenz et al. (2015); Alarcón et al. (2020); Lee et al. (2015); Arrington and Sick (2015); Horbatsch et al. (2017); Lin et al. (2021b, 2022); Cui et al. (2021a); Borah et al. (2020); Borisyuk (2010); Graczyk and Juszczak (2014); Belushkin et al. (2007); Goharipour et al. (2025a). These results cover the interval fm, with a simple combination of fm. This value is smaller than the combined experimental results but larger than the combined theoretical results. It is also worth noting that the difference between the combined experimental and phenomenological results (0.016 fm) is more significant compared to the case of . Finally, combining all experimental, theoretical, and phenomenological results, we obtain fm.